Project

General

Profile

NDOS Commissioning Meeting 2011/10/27

Attendance: Jon, Gavin, Martin, Kurt, Gary, Athans, Mat, Susan, Jan, Zeb, Tom, David, Enrique, Leon, Peter

Agenda:

  1. APD Redistribution Update
  2. DCS Update
  3. DAQ Update
  4. AOB

APD Redistribution Update

  • Mat:
    • finishing redist. today.
    • We will be short one plane (plane 19 in block 2).
    • muon catcher is complete except for 1
    • tracking planes completed
    • dcms 3-2 and 3-3 will be introduced back into the mix.
  • Martin: all set with powering of DCMs? Mat: need to do "setup-online" in order for scripts to work.
  • Jon: how are things going with cooled APDs? Leon: down to 2 of the recovered and cooled APDs; 5 just arrived at CalTech for tests. Parylene-coated APDs are still working; the one that had the known leak started to show some issues, but the problem goes away when it is warmed up.

DCS Update (Martin)

  • Martin: asked Leon if he should change the setpoint on a FEB that has been alarming on and off recently. Leon: probably won't hurt. Martin will try lowering it by half of a degree or so.

DAQ Update

  • Peter:
    • tried to update some software on the NDOS, but ran into a problem where some new ups products were missing. They are being installed now.
    • probably some more updates tomorrow

AOB